Given f(x)= 2/x - 4 and g(x)= +sqrt+1%E2%88%92x2+ determine:
a) f(x+h)−f(x) / 4
(b)g(2−h)

Answer by CubeyThePenguin(3113)

a) f(x+h) - f(x)/4 = 2/(x+h) - 4 - 2/4x - 1 = (2x - 2(x+h)/(4(x)(x+h)) - 5

b) g(2-h) = sqrt(1) = 1